OpenStudy (anonymous):

The angle of elevation from a point on level ground to the top of a tower°4020'. The angle of elevation of the top of the tower from a point 164 feet further back on the same horizontal line is°2210'. Find the height of the tower.

OpenStudy (zpupster):

 tan40=h/x ; tan22.=h/(x+164) tan40x=h tan22(x+164) = h tan40x=tan22x + tann22(164) (tan40 - tan22)x=tan 122(164) solve for x once you have x solve for h using tan40x=h

OpenStudy (anonymous):

Refer to the attached solution from the Mathematica 9 computer program.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your welcome.